How many grams of lithium hydride are necessary to completely react with 19.5 grams of titanium(IV) chloride in the reaction below?
1
Expert's answer
2020-11-12T07:11:31-0500

"\begin{aligned}
4LiH + TiCl_4 \to Ti + 4LiCl + 2H_2
\end{aligned}"1 mole of TICl4​ reacts with 4 moles of LiH


1 mole of "TiCl_4" = 48 + 4(35.5) = 190g

4 moles of "LiH" = 4(7 + 1) = 4(8) = 32g



"\therefore" 190g of "TiCl_4" = 32 g of "LiH"

19.5g of "TiCl_4" = xg


x = "\dfrac{19.5×32}{190} = 3.28g"



"\therefore" 3.28grams of lithium hydride is necessary to completely react with 19.5 grams of titanium(IV) chloride.